The standard quantity and standard price of raw materials required for one unit of product a are given as follows:

Material

Quantity

Selling price

X

2kg

$3 per kg

Y

4kg

$2 per kg

The actual production and relevant data are as follows: Output for product A is 500 units

Material

Total Quantity for 500 units

Total cost ($)

X

1,100kg

3410

Y

1,800kg

3960

Required: Calculate

  1. Material cost variance
  2. Material price variance
  3. Material usage Variance
  4. Material mixture variance
  5. Material Yield variance
Answer #1
Answer
Working
Material SQ AQ
X 1000 1100
Y 2000 1800
Material SP AP
X $         3.0 $         3.1
Y $         2.0 $         2.2
A Material cost variance
Price Variance +Usage Variance
470 A + 100 F
370 A
B Material price variance
(SP-AP)*AQ
X=(3-3.1)*1100 =110 A
Y=(2-2.2)*1800 = 360 A
Total = 470 A
C Material usage Variance
(SQ-AQ)*SP
X=(1000-1100)*3 = 300 A
Y=(2000-1800)*2 = 400 F
Total = 100 F
D Material mixture variance
(SM-AM)*SP
X=(966.67-1100)*3 = 399.99 A
Y=(1933.33-1800)*2      = 266.66 F
Total    = 133.33 A
E

Material Yield variance

( SQ-SM)*SP
X=(1000-966.67)*3 =99.99 F
Y=(2000-1933.33)2 = 133.34 F
Total = 233.33 F
Answer #2
I don't know
source: Answer
answered by: Amita
Answer #3
I don't know
source: Answer
answered by: Amita